OpenStudy (liizzyliizz):

find the derivative of -4e^x at (1,-4e) i just want to know what do you do when u have e..

OpenStudy (anonymous):

\[\frac{d e^x}{dx} = e^x\]

OpenStudy (amistre64):

when you have e you sit back and relax, you let the trancendense of it flow thru you and around you in pulses of yins and yangs lol

OpenStudy (liizzyliizz):

lol if only that gave me an answer, well a correct one. :]

OpenStudy (amistre64):

the derivative of e^x is itself; e^x -4e^x derives to -4e^x

OpenStudy (amistre64):

when x=1; then the function equals -4e

OpenStudy (liizzyliizz):

that's it? lol ugh why do i always over think the simple problems -_-